# Impact of 5G SA Logical Vulnerabilities on UAV Communications: Threat Models and Testbed Evaluation

### 04 Mar 2026 v1

@@ V Conclusion
This work analyzed the impact of logical vulnerabilities in 5G Standalone networks on communication systems between UAVs and control stations. Three threat models were considered, representing distinct attacker positions within the network architecture.
Experiments conducted in the testbed demonstrated that attacks carried out by a malicious device connected to the same slice, by an adversary with access to the network core, or by a compromised gNodeB can directly affect UAV operation. Depending on the scenario, interruption of data sessions, manipulation of navigation commands, and execution of malicious commands were observed.
The results indicate that the security of UAV communication depends on mechanisms distributed across different layers of the system. Network configurations that allow direct communication between devices in the same data domain can facilitate lateral attacks. Internal core interfaces, when not protected by cryptographic mechanisms, can allow manipulation of sessions at the user plane. Application protocols that do not use authentication or integrity also amplify the impact of these attacks.
These observations reinforce the need to consider multi-layered protection in the architecture, including isolation mechanisms in the 5G network and integrity protection in the UAV control protocol.

### 12 Mar 2026 v2

@@ V Conclusion
This work aimed to investigate the impact of logical vulnerabilities in 5G SA networks on UAV operations. When UAV command-and-control (C2) traffic is carried over the 5G User Plane, attacks that target logical components of the network (a rogue UE, the N4 interface, or a compromised gNodeB) can disrupt or manipulate the channel even when the physical radio link remains operational.
To evaluate these scenarios, we implemented a controlled and reproducible 5G testbed that integrates Open5GS, UERANSIM, and MAVLink-based UAV and GCS scripts. The testbed emulates a 3GPP-aligned 5G SA network and carries C2 traffic over the User Plane between UAV and GCS pods. We defined three threat models and implemented representative attacks: (i) C2 injection and GCS spoofing by a rogue UE in the same slice and DNN; (ii) PDU session teardown via PFCP Session Deletion Requests by an insider with N4 access; and (iii) navigation hijacking via GTP-U payload manipulation at a compromised gNodeB. In all models and experiments, we observed a concrete impact on UAV behavior, such as forced landing, failsafe activation, or redirection to attacker-defined coordinates.
Our attacks showed that logical attacks on 5G architecture can compromise UAV C2 without requiring physical access to the radio link or breaking air-interface encryption. The combination of optional or absent integrity protection on N3 and N4 and the lack of application-layer authentication in MAVLink (when message signing is disabled) creates a cross-layer vulnerability that a rogue UE, an insider, or a compromised gNodeB can exploit. This work bridges the gap between telecommunications security and UAV cyber-physical systems by providing testbed evidence and mitigations (MAVLink signing, N4 protection, N3 integrity) that operators and system designers can apply.
Future work includes extending the testbed to evaluate mitigation strategies in a systematic way, testing in real or production-like cellular deployments where radio and core behavior may differ, and scaling the evaluation to multi-UAV or swarm scenarios. The same methodology can be applied to other critical applications that rely on 5G for control or telemetry, and to future 6G architectures.